OpenStudy (anonymous):

Use the three steps to solve the problem. Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture?

OpenStudy (anonymous):

Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture? {maple syrup: gallons, corn syrup: gallons} let quantity of maple syrup used be x gallons corn syrup used will be 50-x gallons 6*x+0,8(50-x)= 2.36*50 6x+40-0.8x=118.00 5.2x=118-40 5,2x=74 x=74/5.2 x=14.2 gallons 35.8 gallons will be corn syrup.